4)A rocket of mass 2*10(4)kg accelerates at 200m/s2 for 30s from an initial speed of 5000m/s
a) How fast will be rocket travelling after the 30s?
b) How much kinetic energy has the rocket gained?
1
Expert's answer
2019-08-07T10:00:04-0400

a) After the 30s the speed of the rocket would be "v=v_0+at=11000\:\mathrm{ m/s}", where "v_0" is the initial speed and "a" is the acceleration.

b) The rocket gained "E=\frac{mv^2}{2}-\frac{mv^2_0}{2}=9.6\cdot 10^{11}\:\mathrm{J}" of kinetic energy within this 30s.